# CRAFT: Characterizing and Root-Causing Fault Injection Threats at Pre-Silicon

### v1

@@ I Introduction
Fault injection attacks (FIAs) deliberately introduce faults into a system to alter its behavior, exploiting, e.g., timing characteristics of designs to induce malfunctions. An attacker may be driven by various goals, such as (a) bypassing security mechanisms to gain unauthorized access to sensitive resources in PIN-based access controls [1]; (b) extracting secret keys by inducing faults during cryptographic operations [2]; (c) disrupting critical system functionality, causing failures in applications ranging from industrial control systems to IoT devices [3]; (d) escalating privileges within embedded systems, enabling unauthorized execution of commands or access to restricted data [4], and (e) targeting data integrity by introducing errors into computation/data flow, compromising the reliability and accuracy of critical systems [5, 6].

### v2

@@ I Introduction
Fault injection attacks (FIAs) deliberately introduce faults into a system to alter its behavior, exploiting, e.g., timing characteristics of designs to induce malfunctions. An attacker may be driven by various goals, such as (a) bypassing security mechanisms to gain unauthorized access to sensitive resources in PIN-based access controls [1, 2, 3]; (b) extracting secret keys by inducing faults during cryptographic operations [4, 5, 6]; (c) disrupting critical system functionality, causing failures in applications ranging from industrial control systems to IoT devices [7, 8, 9, 10, 11, 12, 13, 14]; (d) escalating privileges within embedded systems, enabling unauthorized execution of commands or access to restricted data [15], and (e) targeting data integrity by introducing errors into computation/data flow, compromising the reliability and accuracy of critical systems [16, 17, 18].